OpenStudy (anonymous):

Solving for x: (x^2- a^2)= 3(x+a)

OpenStudy (anonymous):

\[x^2-a^2-3\left( x+a \right)=0\] \[\left( x+a \right)\left( x-a \right)-3\left( x+a \right)=0\] \[\left( x+a \right)\left( x-a-3 \right)=0\] equate each factor to 0 and find x

OpenStudy (anonymous):

How do you get from (x+a)(x-a)-3(x+a)=0 to (x+a)(x-a-3)=0 ??

OpenStudy (anonymous):

\[a^2-b^2=\left( a+b \right)\left( a-b \right)\] i have taken x+a common from both factors.